OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a vertex at the origin and a focus at (0, 9). I need help

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

y=a(x-h)^+k (h,k) is the vertex

OpenStudy (anonymous):

since the vertex is at the origin, it is (0,0)

OpenStudy (anonymous):

wait so what am I suppose to do?

OpenStudy (anonymous):

@abb0t @jhonyy9 @koymoi

OpenStudy (jdoe0001):

so the vertex is at the origin, and the focus is at (0, 9) what's the distance from the vertex to the focus?

OpenStudy (anonymous):

9 right/

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

right so, notice for the focus to the origin, the "x" coordinate didn't change, the "y" did that means is a parabola moving upwards so the "focus form" of a parabola like so is \(\bf (x-h)^2=4p(y-k)\) (h, k) as the vertex p = distance from the vertex to the focus p > 0, opens upward p < 0, opens downward

OpenStudy (anonymous):

ok that makes a little sense

OpenStudy (anonymous):

so am I suppose to use that formula? how do I know what to plug in?

OpenStudy (jdoe0001):

well, your vertex is at the origin, and your focal distance is 9

OpenStudy (anonymous):

I'm not sure how doing that formula will give me one of my answer options, they seem like they don't match

OpenStudy (anonymous):

origin (0,0)

OpenStudy (anonymous):

0+p=9

OpenStudy (anonymous):

now plug them in ;)

OpenStudy (anonymous):

what is p suppose to be?

OpenStudy (jdoe0001):

(h, k) as the vertex p = distance from the vertex to the focus p > 0, opens upward p < 0, opens downward

OpenStudy (anonymous):

so p is 9

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

so what's the next step after finding p?

OpenStudy (jdoe0001):

plug the values in \(\bf (x-h)^2=4p(y-k)\)

OpenStudy (anonymous):

actually if you know the vertex is at its origin, the formula could just be x^=4py or y^=4px

OpenStudy (anonymous):

then like jdoe said you have to decide which one to use based on the value of p

OpenStudy (anonymous):

I am seriously so confused

OpenStudy (anonymous):

so x^=4py you know p=9

OpenStudy (anonymous):

I am just suppose to have an answer in terms of y or y^2 apparently

OpenStudy (anonymous):

plug it in you get x^=36y

OpenStudy (anonymous):

after that what do I do? I mean my answer options are y = 1/36x2 y = 1/9x^2 y^2 = 9x y^2 = 36x

OpenStudy (jdoe0001):

what did you get for the equation?

OpenStudy (anonymous):

for the main one from earlier?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):